# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

### v3

@@ 3.2.3. Splitting Azumaya algebras
${\mathcal{V}}_{(1,\zeta^{\lambda})}={\mathcal{B}}\times\{(1,\zeta^{\lambda})\}={\mathcal{B}}\quad\text{in}\quad{\mathcal{V}}.$
@@
The equivalence (3.3) above depends on the choice of splitting bundles. We will fix and recall the construction of splitting bundles ${\mathfrak{M}}_{\lambda}$ given in [42, §5.4] as follows. Firstly, we consider the completion of ${\mathfrak{U}}_{\zeta}$ at $(1,\chi_{-\rho})\in G^{*}\times_{T/W}T/(W,\bullet)$. By [15, Thms 2.5 and 4.9], the algebra $({\mathfrak{U}}_{\zeta})_{\widehat{(1,\chi_{-\rho})}}$ is a split Azumaya algebra over its center
@@
Hence ${\mathfrak{M}}_{{-\rho}}=\varpi^{*}M_{\chi_{-\rho}}$ is a splitting bundle for $\widetilde{{\mathcal{D}}}_{\zeta}$ on $\mathrm{FN}({{\mathcal{V}}}_{(1,\zeta^{-2\rho})})$. In general, in [43] the author constructed a locally free ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$-module ${\mathcal{O}}^{\lambda}_{{\mathcal{B}}_{\zeta}}$ of rank 1 for any $\lambda\in\Lambda$. By [42, Thm 5.3], one can set ${\mathfrak{M}}_{\lambda}={\mathcal{O}}_{{\mathcal{B}}_{\zeta}}^{\lambda+\rho}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathfrak{M}}_{{-\rho}}$.
@@ 3.2.4. Formal neighborhood of ${\mathcal{B}}$
$\mathrm{FN}_{{\mathcal{V}}}({\mathcal{B}}\times\{(1,\zeta^{\lambda})\})=\mathrm{FN}_{\widetilde{G^{*}}}({\mathcal{B}})=\mathrm{FN}_{\widetilde{G}}({\mathcal{B}}),\quad\forall\lambda\in\Lambda.$
@@
${\mathfrak{g}}_{\hat{0}}{\ \xrightarrow{\sim}\ }G_{\hat{1}}\quad\text{and}\quad\tilde{{\mathfrak{g}}}_{\hat{0}}{\ \xrightarrow{\sim}\ }\widetilde{G}_{\hat{1}}.$
Hence we can identify $\mathrm{FN}_{\widetilde{G}}({\mathcal{B}})=\mathrm{FN}_{\tilde{{\mathfrak{g}}}}({\mathcal{B}})$, and it follows that
$\mathrm{Coh}_{{\mathcal{B}}\times\{(1,\zeta^{2\lambda})\}}({\mathcal{V}})=\mathrm{Coh}_{{\mathcal{B}}}(\tilde{{\mathfrak{g}}}),\quad\forall\lambda\in\Lambda.$
@@
(3.6) $D^{b}\mathrm{Coh}_{{\mathcal{B}}}(\tilde{{\mathfrak{g}}}){\ \xrightarrow{\sim}\ }D^{b}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(1,\chi_{0})}).$
We set ${\mathrm{A}}_{0}$ be the specialization of ${\mathrm{A}}$ at $0\in{\mathfrak{g}}$.
@@
The equivalence (3.6) is exact for the exotic t-structure on the LHS and the trivial t-structure on the RHS. Therefore, there are equivalences
(3.7) ${\mathrm{A}}\text{-}\mathrm{mod}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(1,\chi_{0})}$
@@
${\mathrm{A}}_{0}\text{-}\mathrm{mod}{\ \xrightarrow{\sim}\ }u_{\zeta}\text{-}\mathrm{mod}_{\chi_{0}}.$
@@ 3.2.5. Graded version
By [42, §7.4] there exists a $\Lambda$-grading (in completion sense) on $M_{\chi_{-\rho}}$. It yields an equivalence
@@
Moreover, the splitting bundles ${\mathfrak{M}}_{\lambda}$, $\lambda\in\Lambda$ are endowed with $T$-equivariant structures (in completion sense) by their constructions in §3.2.3.
We have the following graded version of Proposition 3.5.
@@
The $T$-equivariant structure on ${\mathfrak{M}}_{{0}}$ induces equivalences
(3.9) ${\mathrm{A}}\text{-}\mathrm{mod}^{T}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})},$
@@
${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{{\mathbb{C}}}.$
@@ 4. Equivalence ${\mathfrak{F}}^{b}$
${\mathfrak{b}}_{S}={\mathfrak{b}}\times_{\mathfrak{t}}{\mathfrak{t}}_{\hat{0}},\quad{\mathrm{A}}^{b}={\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{b}}],\quad\text{and}\quad{\mathrm{A}}^{b}_{S}={\mathrm{A}}^{b}\otimes_{{\mathbb{C}}[{\mathfrak{b}}]}{\mathbb{C}}[{\mathfrak{b}}_{S}].$
The exponential map induces isomorphisms of schemes ${\mathfrak{n}}{\ \xrightarrow{\sim}\ }N$ and ${\mathfrak{b}}_{S}{\ \xrightarrow{\sim}\ }B\times_{T}T_{\hat{1}}$. Since the equivalence (3.9) is ${\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}$-linear, it specializes to an equivalence
(4.1) ${\mathfrak{F}}^{\prime}:\ {\mathrm{A}}^{b}\text{-}\mathrm{mod}^{T}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})}.$
@@
${\mathfrak{F}}^{b}:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$
@@
4.1. Equivalence
Note that ${\mathrm{A}}^{b}\text{-}\mathrm{mod}^{T}_{0}$ is the full subcategory of ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ of the modules set-theoretically supported at $0\in{\mathfrak{b}}_{S}$. We also view ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})}$ as a full subcategory of ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ as follows. Let $M$ be in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})}$, then by definition the operator $\zeta^{-(\mu,\lambda)}K_{\mu}$ on $M_{\lambda}$ is unipotent for any $\mu,\lambda\in\Lambda$. Hence the assignment
@@ 4.1. Equivalence
Since the reflection functors $\Theta^{c}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$) on ${\mathrm{A}}\text{-}\mathrm{mod}$ are $G$-equivariant and ${\mathbb{C}}[{\mathfrak{g}}]$-linear, they induce reflection functors (still denoted by $\Theta^{c}_{s}$) on ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{K}$ by base change along ${\mathfrak{b}}_{S}\rightarrow{\mathfrak{g}}$, for any subgroup $K$ in $B$.
@@
(4.2) ${\mathfrak{F}}^{b}:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S},$
@@
$S_{n}=S/({\mathfrak{m}}_{1}\cap{\mathbb{C}}[T])^{n+1}S,\qquad{\mathrm{A}}^{b}_{n}={\mathrm{A}}^{b}_{S}/{\mathfrak{m}}_{0}^{n+1}{\mathrm{A}}^{b}_{S}\quad\text{and}\quad{\mathfrak{U}}^{b}_{\zeta,n}={\mathfrak{U}}^{b}_{\zeta}/{\mathfrak{m}}_{1}^{n+1}{\mathfrak{U}}^{b}_{\zeta}.$
Since ${\mathfrak{F}}^{\prime}$ is ${\mathbb{C}}[{\mathfrak{b}}]_{\hat{0}}$-linear, it yields equivalences
${\mathfrak{F}}^{\prime}:\ {\mathrm{A}}^{b}_{n}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}},\quad\forall n\geq 0.$
@@
$\tau_{n}:{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}\rightarrow{\mathrm{A}}^{b}_{n}\text{-}\mathrm{mod}^{T},\quad\tau_{n}:{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}\rightarrow{\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}},$
and ${\mathfrak{F}}^{\prime}$ is clearly compatible with $\tau_{n}$. Note that ${\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{n})$ is projective in ${\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}}$, the family $\{{\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{n})\}_{n}$ admits a lifting $P^{b}$ in $\mathrm{Proj}({\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S})$. We choose a family of isomorphisms ${\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{n})\cong\tau_{n}(P^{b})$, $n\geq 0$ that are compatible with specializations. To proceed, we need the following lemmas.
@@
(1) Since $\tilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$ is a projective morphism to an affine scheme, ${\mathrm{A}}=\mathrm{End}^{\mathrm{op}}_{\tilde{{\mathfrak{g}}}}({\mathcal{E}})$ is a finitely generated ${\mathbb{C}}[{\mathfrak{g}}]$-module by Serre’s Theorem. Hence ${\mathrm{A}}^{b}_{S}$ is a finitely generated ${\mathbb{C}}[{\mathfrak{b}}_{S}]$-module. Note that ${\mathrm{A}}^{b}_{S}$ is moreover $\Lambda$-graded and complete over $S$. We view ${\mathbb{C}}[{\mathfrak{b}}_{S}]$ as a ${\mathbb{Z}}_{\geq 0}$-graded ring via ${\mathbb{C}}[{\mathfrak{b}}_{S}]_{i}=\bigoplus_{\mathrm{ht}(\eta)=i}{\mathbb{C}}[{\mathfrak{b}}_{S}]_{-\eta}$, and consider the graded ideal ${\mathfrak{m}}_{0}{\mathbb{C}}[{\mathfrak{b}}_{S}]$, then our assertion follows from Lemma 4.2.
@@
where the first and the second last equalities are by Lemma 4.3. Note that the simple modules of ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ are contained in $u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}={\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{0}\text{-}\mathrm{mod}^{T})$, and therefore they admit a surjection from $\bigoplus_{\nu}{\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{1}\otimes{\mathbb{C}}_{\nu})=\bigoplus_{\nu}\tau_{1}(P^{b}\otimes{\mathbb{C}}_{l\nu})$. It implies that $\{P^{b}\otimes{\mathbb{C}}_{l\nu}\}_{\nu}$ forms a generating family of projective modules in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$. Hence there is a Morita equivalence from ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ to the category of $\Lambda$-graded modules of the algebra $\bigoplus_{\nu}\mathrm{Hom}\big{(}P^{b},P^{b}\otimes{\mathbb{C}}_{l\nu})$. Now our desired equivalence (4.2) follows from the isomorphism (4.4).
By Proposition 3.6, the equivalence (4.1) is compatible with reflection functors and $\Lambda$-translations. The compatibility for (4.2) follows. ∎
We set ${\mathrm{A}}_{\mathfrak{n}}:={\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]$.
@@
(4.5) ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}.$
4.2. Steinberg block and functor ${\mathbb{V}}^{b}$
4.2.1. Equivalence for Steinberg block
Recall the equivalence (3.4). By the ${\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}$-linearity it specializes to an equivalence

### v4

@@ 3.2.3. Splitting Azumaya algebras
${\mathcal{V}}_{(1,\zeta^{\lambda})}={\mathcal{B}}\times\{(1,\zeta^{\lambda})\}\simeq{\mathcal{B}}\quad\text{in}\quad{\mathcal{V}}.$
@@
The equivalence (3.3) above depends on the choice of splitting bundles. We will fix and recall the construction of splitting bundles ${\mathfrak{M}}_{\lambda}$ given in [39, §5.4] as follows. Firstly, we consider the completion of ${\mathfrak{U}}_{\zeta}$ at $(1,\chi_{-\rho})\in G^{*}\times_{T/W}T/(W,\bullet)$. By [15, Thms 2.5 and 4.9], the algebra $({\mathfrak{U}}_{\zeta})_{\widehat{(1,\chi_{-\rho})}}$ is a split Azumaya algebra over its center
@@
Hence ${\mathfrak{M}}_{{-\rho}}=\varpi^{*}M_{\chi_{-\rho}}$ is a splitting bundle for $\widetilde{{\mathcal{D}}}_{\zeta}$ on $\mathrm{FN}({{\mathcal{V}}}_{(1,\zeta^{-2\rho})})$. In general, in [41] the author constructed a locally free ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$-module ${\mathcal{O}}^{\lambda}_{{\mathcal{B}}_{\zeta}}$ of rank 1 for any $\lambda\in\Lambda$. By [39, Thm 5.3], one can set ${\mathfrak{M}}_{\lambda}={\mathcal{O}}_{{\mathcal{B}}_{\zeta}}^{\lambda+\rho}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathfrak{M}}_{{-\rho}}$.
@@ 3.2.4. Formal neighborhood of $\mathcal{B}$}
$\mathrm{FN}_{{\mathcal{V}}}({\mathcal{B}}\times\{(1,\zeta^{\lambda})\})\simeq\mathrm{FN}_{\widetilde{G^{*}}}({\mathcal{B}})\simeq\mathrm{FN}_{\widetilde{G}}({\mathcal{B}}),\quad\forall\lambda\in\Lambda.$
@@
${\mathfrak{g}}_{\hat{0}}{\ \xrightarrow{\sim}\ }G_{\hat{1}}\quad\text{and}\quad\widetilde{{\mathfrak{g}}}_{\hat{0}}{\ \xrightarrow{\sim}\ }\widetilde{G}_{\hat{1}}$.
[…]
$\mathrm{Coh}_{{\mathcal{B}}\times\{(1,\zeta^{2\lambda})\}}({\mathcal{V}})\simeq\mathrm{Coh}_{{\mathcal{B}}}(\widetilde{{\mathfrak{g}}}),\quad\forall\lambda\in\Lambda.$
@@
(3.6) $D^{\mathrm{b}}\mathrm{Coh}_{{\mathcal{B}}}(\widetilde{{\mathfrak{g}}}){\ \xrightarrow{\sim}\ }D^{\mathrm{b}}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(1,\chi_{0})}).$
We set ${\mathsf{A}}_{0}$ be the specialization of $\mathsf{A}$} at $0\in{\mathfrak{g}}$.
@@
The equivalence (3.6) is exact for the exotic t-structure on the LHS and the trivial t-structure on the RHS. Therefore, there is an equivalence
(3.7) ${\mathsf{A}}\text{-}\mathrm{mod}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(1,\chi_{0})}.$
@@
${\mathsf{A}}_{0}\text{-}\mathrm{mod}{\ \xrightarrow{\sim}\ }u_{\zeta}\text{-}\mathrm{mod}_{\chi_{0}}.$
@@ 3.2.5. Graded version
By [39, §7.4] there exists a $\Lambda$-grading (in completion sense) on $M_{\chi_{-\rho}}$. It yields an equivalence
@@
Moreover, the splitting bundles ${\mathfrak{M}}_{\lambda}$, $\lambda\in\Lambda$ are thus endowed with $\Lambda$-grading (in completion sense) by their constructions in §3.2.3. We have the following graded version of Proposition 3.5.
@@
The $\Lambda$-grading structure on ${\mathfrak{M}}_{{0}}$ induces an equivalence
(3.9) ${\mathsf{A}}\text{-}\mathrm{mod}^{T}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})},$
@@
${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{{\mathbb{C}}}.$
@@ 4. Equivalence ${\mathfrak{F}}^{b}$
${\mathfrak{b}}_{S}={\mathfrak{b}}\times_{\mathfrak{t}}{\mathfrak{t}}_{\hat{0}},\quad{\mathsf{A}}^{b}={\mathsf{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{b}}],\quad\text{and}\quad{\mathsf{A}}^{b}_{S}={\mathsf{A}}^{b}\otimes_{{\mathbb{C}}[{\mathfrak{b}}]}{\mathbb{C}}[{\mathfrak{b}}_{S}].$
The exponential map induces isomorphisms of schemes ${\mathfrak{n}}{\ \xrightarrow{\sim}\ }N$ and $^{\mathfrak{b}}_{S}{\ \xrightarrow{\sim}\ }B\times_{T}T_{\hat{1}}$. Since the equivalence (3.9) is $^{\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}$-linear, it specializes to an equivalence
(4.1) ${\mathfrak{F}}^{\prime}:\ {\mathsf{A}}^{b}\text{-}\mathrm{mod}^{T}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})}.$
@@
${\mathfrak{F}}^{b}:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$
@@
4.1. Equivalence ${\mathfrak{F}}^{b}$
Note that ${\mathsf{A}}^{b}\text{-}\mathrm{mod}^{T}_{0}$ is the full subcategory of ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ of the modules set-theoretically supported at $0\in{\mathfrak{b}}_{S}$. We also view ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})}$ as a full subcategory of ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ as follows. Let $M$ be in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})}$, then by definition the operator $\zeta^{-(\mu,\lambda)}K_{\mu}$ on $M_{\lambda}$ is unipotent for any $\mu,\lambda\in\Lambda$. Hence the assignment
@@ 4.1. Equivalence ${\mathfrak{F}}^{b}$
Since the reflection functors $\Theta^{c}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$) on ${\mathsf{A}}\text{-}\mathrm{mod}$ are $G$-equivariant and ${\mathbb{C}}[{\mathfrak{g}}]$-linear, they induce reflection functors (still denoted by $\Theta^{c}_{s}$) on ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{K}$ by base change along ${\mathfrak{b}}_{S}\rightarrow{\mathfrak{g}}$, for any subgroup scheme $K$ in $B$.
@@
(4.2) ${\mathfrak{F}}^{b}:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S},$
@@
$S_{n}=S/({\mathfrak{m}}_{1}\cap{\mathbb{C}}[T])^{n+1}S,\qquad{\mathsf{A}}^{b}_{n}={\mathsf{A}}^{b}_{S}/{\mathfrak{m}}_{0}^{n+1}{\mathsf{A}}^{b}_{S}\quad\text{and}\quad{\mathfrak{U}}^{b}_{\zeta,n}={\mathfrak{U}}^{b}_{\zeta}/{\mathfrak{m}}_{1}^{n+1}{\mathfrak{U}}^{b}_{\zeta}.$
Since $\mathfrak{F}$$^{\prime}$ is $^{\mathbb{C}}[{\mathfrak{b}}]_{\hat{0}}$-linear, it yields equivalences
${\mathfrak{F}}^{\prime}:\ {\mathsf{A}}^{b}_{n}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}},\quad\forall n\geq 0.$
@@
$\tau_{n}:{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}\rightarrow{\mathsf{A}}^{b}_{n}\text{-}\mathrm{mod}^{T},\quad\tau_{n}:{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}\rightarrow{\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}},$
and $\mathfrak{F}$$^{\prime}$ is clearly compatible with $\tau_{n}$. Note that ${\mathfrak{F}}^{\prime}({\mathsf{A}}^{b}_{n})$ is projective in ${\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}}$, the family $\{{\mathfrak{F}}^{\prime}({\mathsf{A}}^{b}_{n})\}_{n}$ admits a lifting $P^{b}$ in $\mathrm{Proj}({\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S})$. We choose a family of isomorphisms ${\mathfrak{F}}^{\prime}({\mathsf{A}}^{b}_{n})\cong\tau_{n}(P^{b})$, $n\geq 0$ that are compatible with specializations. To proceed, we need the following lemmas.
@@
(1) Since $\widetilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$ is a projective morphism to an affine scheme, ${\mathsf{A}}=\mathrm{End}^{\mathrm{op}}_{\widetilde{{\mathfrak{g}}}}({\mathcal{E}})$ is a finitely generated ${\mathbb{C}}[{\mathfrak{g}}]$-module by Serre’s Theorem. Hence ${\mathsf{A}}^{b}_{S}$ is a finitely generated ${\mathbb{C}}[{\mathfrak{b}}_{S}]$-module. Note that ${\mathsf{A}}^{b}_{S}$ is moreover $\Lambda$-graded and complete over $S$. We view ${\mathbb{C}}[{\mathfrak{b}}_{S}]$ as a ${\mathbb{Z}}_{\geq 0}$-graded ring via ${\mathbb{C}}[{\mathfrak{b}}_{S}]_{i}=\bigoplus_{\mathrm{ht}(\eta)=i}{\mathbb{C}}[{\mathfrak{b}}_{S}]_{-\eta}$, and consider the graded ideal ${\mathfrak{m}}_{0}{\mathbb{C}}[{\mathfrak{b}}_{S}]$, then our assertion follows from Lemma 4.2.
@@
where the first and the second last equalities are by Lemma 4.3. Note that the simple modules of ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ are contained in $u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}={\mathfrak{F}}^{\prime}({\mathsf{A}}^{b}_{0}\text{-}\mathrm{mod}^{T})$, and therefore they admit a surjection from $\bigoplus_{\nu}{\mathfrak{F}}^{\prime}({\mathsf{A}}^{b}_{1}\otimes{\mathbb{C}}_{\nu})=\bigoplus_{\nu}\tau_{1}(P^{b}\otimes{\mathbb{C}}_{l\nu})$. It implies that $\{P^{b}\otimes{\mathbb{C}}_{l\nu}\}_{\nu}$ forms a generating family of projective modules in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$. Hence there is a Morita equivalence from ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ to the category of $\Lambda$-graded modules of the algebra $\bigoplus_{\nu}\mathrm{Hom}\big(P^{b},P^{b}\otimes{\mathbb{C}}_{l\nu})$. Now our desired equivalence (4.2) follows from the isomorphism (4.4).
Finally, by Proposition 3.6, the equivalence (4.1) is compatible with reflection functors and $\Lambda$-translations. The compatibility for (4.2) follows. ∎
We set ${\mathsf{A}}_{\mathfrak{n}}:={\mathsf{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]$.
@@
(4.5) ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}.$
4.2. Equivalence ${\mathfrak{G}}^{b}$ for Steinberg block
Recall the equivalence (3.4). By the $^{\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}$-linearity it specializes to an equivalence